OpenStudy (anonymous):

I really need help what are all the zeros for y=x(x+5)(x+4)

OpenStudy (kc_kennylau):

0,-5,-4

OpenStudy (kc_kennylau):

x(x+5)(x+4)=0 x=0 or x+5=0 or x+4=0

OpenStudy (anonymous):

thanks

OpenStudy (kc_kennylau):

no problem

OpenStudy (akashdeepdeb):

y=x(x+5)(x+4) To find the zeroes equate y to 0 then y=x(x+5)(x+4) = 0 Thus First zero can be =>x = 0/(x+5)(x+4) = 0 Second zero can be => x+4 = 0/x(x+5) => x+4=0 => x = -4 Third zero can be => x+5 = 0/x(x+4) => x+5 =0 => x =-5 And thus the three zeroes of this third degree polynomial will be 0,-4,-5 Understood? :)

OpenStudy (phi):

you want the x values that make y = 0 The main idea is if you multiply three numbers, say A * B *C if any is 0, you will get 0. in your case, the three "numbers" are x * (x+5) * (x+4) if any of the three are zero, that is an answer. you get 3 different values